# The Adaptivity Barrier in Batched Nonparametric Bandits: Sharp Characterization of the Price of Unknown Margin

### v1

@@ A.3 Proof of the upper bound
$R_{T}(\hat{\pi},\mathcal{P})$ $\apprle t_{1}+\sum_{i=2}^{M-1}(t_{i}-t_{i-1})\cdot w_{i-1}^{\beta+\alpha\beta}+(T-t_{M-1})w_{M-1}^{\beta+\alpha\beta}.$
@@ A.4 Proof of the lower bound
The proof mainly follows the strategy outlined in [22], but with a slightly different reward function construction to handle the wider range of $\alpha$.
@@ A.4.1 Construction of the hard instances
$\xi_{m,j}(x):=\delta_{m}\,\phi^{\beta}\big(Mz_{m}\|x-q_{m,j}\|_{\infty}\big)\,\mathbf{1}{\{x\in C_{m,j}\}},\qquad\delta_{m}:=D_{\phi}\,(Mz_{m})^{-\beta},$ (31)
@@
$|S_{m}|=s_{m}:=\Big\lceil M^{-1}(Mz_{m})^{\,d-\alpha\beta}\Big\rceil,\quad 2\leq m\leq M,\quad|S_{1}|=s_{1}:=M^{d-1}.$ (32)
@@ A.4.2 Lower bounding the regret during the $m$-th batch
Appendix B On the variational problem
@@
Define $u_{0}:=0$ and $u_{M}:=1$. For $i=1,\dots,M$, set
$\bm{d}_{i}\;:=\;\bm{e}_{i-1}-\bm{e}_{i}\in\mathbb{R}^{M-1},\qquad\text{with the convention }\bm{e}_{0}:=\bm{0},\ \bm{e}_{M}:=\bm{0}.$
@@ B.2.1 Monotonicity of $\{\gamma_{i}\}$
$v_{L}:=-h_{M}^{\prime}(\gamma_{\mathrm{max}}),\quad v_{U}:=-h_{M}^{\prime}(\gamma_{\mathrm{min}})\quad\text{with }0<v_{L}<v_{U}.$
@@ Appendix C Remaining proofs for the lower bound
$r(\sigma):=\frac{(\tfrac{1}{2}+\sigma\delta)^{R}(\tfrac{1}{2}-\sigma\delta)^{N-R}}{(\tfrac{1}{2})^{N}}=(1+2\sigma\delta)^{R}(1-2\sigma\delta)^{N-R},\quad\sigma\in\{+1,-1\},$
@@
where the last step applies Lemma 1.1 in [6]. Using the PGF of binomial distribution, we obtain
@@ Appendix D Proof of the upper bound
This section is devoted to establishing Lemma 1, whose proof follows the framework developed in [31, 22].
@@ D.1 Introducing the good events
Define the event $\mathcal{A}_{C}=\{\underline{\mathcal{I}}_{C}\subseteq\mathcal{I}_{C}^{\prime}\subseteq\bar{\mathcal{I}}_{C}\}$. Besides, define $\mathcal{G}_{C}=\cap_{C^{\prime}\in\mathcal{P}(C)}\mathcal{A}_{C^{\prime}}$. For $i\geq 1$, recall that $\mathcal{B}_{i}$ is the collection of bins $C$ with $|C|=(\prod_{l=0}^{i-1}g_{l})^{-1}=w_{i}$. The following lemma adapted from [22] shows successive elimination succeeds with high probability.

### v2

@@ A.3 Proof of the upper bound
$R_{T}(\hat{\Gamma},\hat{\pi};P)$ $\apprle t_{1}+\sum_{i=2}^{M-1}(t_{i}-t_{i-1})\cdot w_{i-1}^{\beta+\alpha\beta}+(T-t_{M-1})w_{M-1}^{\beta+\alpha\beta}.$
@@ A.4 Proof of the lower bound
The proof mainly follows the strategy outlined in Jiang and Ma (2025), but with a slightly different reward function construction to handle the wider range of $\alpha$.
@@ A.4.1 Construction of the hard instances
$\xi_{m,j}(x)\coloneqq\delta_{m}\,\phi^{\beta}\big(Mz_{m}\|x-q_{m,j}\|_{\infty}\big)\,\mathbf{1}{\{x\in C_{m,j}\}},\qquad\delta_{m}\coloneqq D_{\phi}\,(Mz_{m})^{-\beta},$ (31)
@@
$|S_{m}|=s_{m}\coloneqq\Big\lceil M^{-1}(Mz_{m})^{\,d-\alpha\beta}\Big\rceil,\quad 2\leq m\leq M,\quad|S_{1}|=s_{1}\coloneqq M^{d-1}.$ (32)
@@ A.4.2 Lower bounding the regret during the $m$-th batch
Appendix B Detailed analysis of the variational problem
@@
Define $u_{0}\coloneqq 0$ and $u_{M}\coloneqq 1$. For $i=1,\dots,M$, set
$\bm{d}_{i}\;\coloneqq\;\bm{e}_{i-1}-\bm{e}_{i}\in\mathbb{R}^{M-1},\qquad\text{with the convention }\bm{e}_{0}\coloneqq\bm{0},\ \bm{e}_{M}\coloneqq\bm{0}.$
@@ B.2.1 Monotonicity of $\{\gamma_{i}\}$
$v_{L}\coloneqq-h_{M}^{\prime}(\gamma_{\mathrm{max}}),\quad v_{U}\coloneqq-h_{M}^{\prime}(\gamma_{\mathrm{min}})\quad\text{with }0<v_{L}<v_{U}.$
@@ Appendix C Remaining proofs for the lower bound
$r(\sigma)\coloneqq\frac{(\tfrac{1}{2}+\sigma\delta)^{R}(\tfrac{1}{2}-\sigma\delta)^{N-R}}{(\tfrac{1}{2})^{N}}=(1+2\sigma\delta)^{R}(1-2\sigma\delta)^{N-R},\quad\sigma\in\{+1,-1\},$
@@
where the last step applies Lemma 1.1 in Bardenet and Maillard (2015). Using the PGF of binomial distribution, we obtain
@@ Appendix D Proof of the upper bound
This section is devoted to establishing Lemma 1, whose proof follows the framework developed in Perchet and Rigollet (2013); Jiang and Ma (2025).
@@ D.1 Introducing the good events
Define the event $\mathcal{A}_{C}=\{\underline{\mathcal{I}}_{C}\subseteq\mathcal{I}_{C}^{\prime}\subseteq\bar{\mathcal{I}}_{C}\}$. Besides, define $\mathcal{G}_{C}=\cap_{C^{\prime}\in\mathcal{P}(C)}\mathcal{A}_{C^{\prime}}$. For $i\geq 1$, recall that $\mathcal{B}_{i}$ is the collection of bins $C$ with $|C|=(\prod_{l=0}^{i-1}g_{l})^{-1}=w_{i}$. The following lemma adapted from Jiang and Ma (2025) shows successive elimination succeeds with high probability.